Hiya guys i need help on these problems:
its like greek to me...dont understand it a bit!
>
The hiking club hiked to Robbers Cave State Park at 4 mph. They got a ride back to town in a truck that went 20 mph. If roundtrip took 18 hours how far was it from town to park?
>
Simon set out on a hike. After walking for some time at 5kph he caught a ride back home in a truck that traveled at 20 kph. If the roundtrip took 10 hours how far did he walk?
and last one...
There were 6 more girls than twice the number of boys. There were 36 boys and girls in all. How many boys and how many were there?

Answer by jai_kos(139)

Solution:
a) The hiking club hiked to Robbers Cave State Park at 4 mph
They got a ride back to town in a truck that went 20 mph.
Formula: Total speed = total distance / total time.
Total speed = (4 + 20) mph = 24mph
Total time = 18hrs
Substituting in the above formula, we have
24 = total distance / 18
24 * 18 = total distance
432 = total distance
Distance from park to the town + distance from the town to park = 432
Distance from park to the town = distance from the town to park
So we write it as,
2 distance = 432
Distance = 432 /2
Distance = 216miles
The distance from the park to the town is 216 miles.
b) we use the similar concept to solve.
Formula: Total speed = total distance / total time.
5 + 20 = total distance / 10
25 = total distance / 10
25 * 10 = total distance
250 = total distance
Distance walked + distance came back to home = 250
Distance walked = distance came back to home
So we write it as,
2 Distance walked = 250
Distance walked = 250 /2
Distance walked = 125 km
c) Let the number of girls be “x”
And the number of boys be “y”
Given that there are 6 more girls than twice the number of boys
i.e. 6x = 2y
3x = y --- (1)
Also given there are 36, boys and girls in all.
i.e. x + y = 36 --- (2)
Substituting equation (1) in equation (2), we get
x + 3x = 36
4x = 36
x = 9
Now substitute the value of “x” in equation (1) we get
3* 9 = y
27 = y
Therefore the number of girls are 9 and number of boys are 27